A quadratic relation has the equation y=a(x-s)(x-t)
 Find the value of a when
 a) y=a(x-2)(x+6) and (3,5) is a point on the graph
 b)the parabola has zeros of 4 and -2 and a y-intercept of 1
 c)the parabola has x-intercepts of 4 and -2 and a y-intercept of -1
 d)the parabola has zeros of 5 and 0 and a minimum value of -10
 e)the parabola has x-intercepts of 5 and -3 and a maximum value of 6
 
 Answer by stanbon(75887)
       
 A quadratic relation has the equation y=a(x-s)(x-t) Find the value of a when
 a) y=a(x-2)(x+6) and (3,5) is a point on the graph
 5 = a(3-2)(3+6)
 5 = a*1*9
 a = 5/9
 ------------------------
 b)the parabola has zeros of 4 and -2 and a y-intercept of 1
 y = a(x-4)(x+2)
 1 = a(-4)(2)
 a = -1/8
 -----------------------
 c)the parabola has x-intercepts of 4 and -2 and a y-intercept of -1
 y = a(x-4)(x+2)
 -1 = a(-4)(2)
 a = 1/8
 ------------------------
 d)the parabola has zeros of 5 and 0 and a minimum value of -10
 y = a(x-5)(x)
 Note: The vertex is midway between 5 and 0: i.e. x = 5/2
 10 = a(5/2 -5)(5/2)
 10 = a(-5/2)(5/2)
 10 = a(-25/4)
 a = -40/25
 a = -8/5
 ---------------------
 e)the parabola has x-intercepts of 5 and -3 and a maximum value of 6
 y = a(x-5)(x+3)
 Max at (1,6)
 6 = a(-4)(4)
 a = -6/16 = -3/8
 ---------------------
 cheers,
 Stan H.
 -------------
 
 | 
  
 | 
 |